what is the answers and steps to this equation? 3(q-5)=2(q+5)

Answer by oscargut(2103)

3(q-5)=2(q+5)
3q-15 = 2q+10
3q-2q = 10+15
q = 25
I can help you with more questions at: mthman@gmail.com
Thanks